OpenStudy (anonymous):

Find the derivative of f(x) = -6/x at x = 12.

OpenStudy (abb0t):

\(f(12)=\large \frac{6}{x^2}\)

OpenStudy (abb0t):

You need to learn your derivative rules.

OpenStudy (anonymous):

I know them, but I'm using them with limits. how you solve them is different @abb0t

OpenStudy (abb0t):

It's just tedious algebra.

OpenStudy (anonymous):

then let me ask, How did you get the x^2?

OpenStudy (anonymous):

@abb0t

OpenStudy (abb0t):

\(\large \frac{1}{x^2} = x^{-2}\)

OpenStudy (anonymous):

But, my answer choices are 24 1/2 1/24 2 @abb0t